OpenStudy (anonymous):

((looking to understand this, not get a direct answer!)) What is the solution of the system? 2x-y+-7 4x-y=-4 A. (-1.5,4) B. (1.5, 10) C. (4, -1.5) D. (-1.5,-2) ((any help is appreciated, thank you!))

OpenStudy (dtan5457):

i would multiply the first equation by -2 so you add them both up and use elimination to solve. 2x-y=-7 4x-y=-4 -4x+2y=14 4x-y=-4

OpenStudy (dtan5457):

what do you get for y?

OpenStudy (dtan5457):

wow once you get y you don't even need to solve for x

OpenStudy (anonymous):

I don't know what y is

OpenStudy (dtan5457):

-4x+2y=14 4x-y=-4 all you have to do is add these 2 up like its normal addition

OpenStudy (dtan5457):

get it?

OpenStudy (anonymous):

getting it so far, yes

OpenStudy (dtan5457):

so add them up to get y lol

OpenStudy (dtan5457):

-4x+2y=14 4x-y=-4 -4x+4x=0 2y+(-y)=y 14+(-4)=10 therefore y=?

OpenStudy (anonymous):

hold on hold on lemme solve this. I'm slow with math, pft. Gimme a bit.

OpenStudy (dtan5457):

haha it's really easy i think you just don't see it

OpenStudy (anonymous):

yeah, I'm probably just not getting it at all...

OpenStudy (dtan5457):

to solve these problems you need to isolate a variable. that's why i multiplied the first one by -2 so that the -4x and 4x will cross out once you add it

OpenStudy (anonymous):

I still don't get it. Sorry.

OpenStudy (dtan5457):

uhh

OpenStudy (dtan5457):

2x-y+=7 4x-y=-4 the absolute easiest way your gonna solve this is to plug these both into the calculator and see where the 2 lines intersect

OpenStudy (dtan5457):

A. (-1.5,4) B. (1.5, 10) C. (4, -1.5) D. (-1.5,-2) or you can plug these points in one by one to 2x-y+=7 4x-y=-4 with the appropriate x and y coordinates

OpenStudy (dtan5457):

personally i think algebraic elimination or substitution is the easiest which is what I tried to show you at first. what are you not understanding?

OpenStudy (anonymous):

Alright I have a calculator now. I know it's not A, so far.

OpenStudy (dtan5457):

alright, solve it with a calculator for now. but do attempt to learn the algebraic methods eventually as some tests may not allow a calculator .

OpenStudy (anonymous):

True true. Eliminated B as well. Thank you for this help.

OpenStudy (anonymous):

I got C.

OpenStudy (dtan5457):

actually it's not C lol

OpenStudy (anonymous):

Alright then, I'll redo my math for the other three :/

OpenStudy (dtan5457):

what does the intersection say lol

OpenStudy (anonymous):

... B

OpenStudy (dtan5457):

yep

OpenStudy (anonymous):

I wasn't using a graphing calculator I don't have one, so I was just using a normal one

OpenStudy (anonymous):

Thank you for the help, haha... Sorry for going in circles with this problem.

OpenStudy (dtan5457):

np. good effort